OpenStudy (anonymous):

Can someone tell me how to solve equations like these? If f(x) = 2x + 1 and g(x) = −2x − 3, find (f + g)(−1). Thank you :)

jimthompson5910 (jim_thompson5910):

(f+g)(x) = f(x) + g(x) (f+g)(x) = 2x + 1 + (-2x - 3) (f+g)(x) = 2x + 1 - 2x - 3 (f+g)(x) = -2

jimthompson5910 (jim_thompson5910):

Now plug in x = -1

OpenStudy (anonymous):

The problem is I'm not understanding the first steps.

jimthompson5910 (jim_thompson5910):

which steps

jimthompson5910 (jim_thompson5910):

step 1? step 2? or something else

OpenStudy (anonymous):

When given an equation like this, If f(x) = 2x + 1 and g(x) = −2x − 3, find (f + g)(−1). What does it mean: Find(f + g)(-1) I'm confused I also see stuff like: Find g(f[4]) And it's really confusing.

jimthompson5910 (jim_thompson5910):

(f + g)(-1) really means f(-1) + g(-1)

jimthompson5910 (jim_thompson5910):

so (f + g)(x) turns into f(x) + g(x)

OpenStudy (anonymous):

So it's saying: What is f and g times -1? And g(f[-1])] is saying, what is f times -1 Then we have: g(f), and we solve that and we've solved for g(f[-])?

jimthompson5910 (jim_thompson5910):

no, g(f(-1)) is a function composition

jimthompson5910 (jim_thompson5910):

you basically evaluate f(-1) first then you take the result of that and plug it into g(x)

OpenStudy (anonymous):

ok thanks

jimthompson5910 (jim_thompson5910):

np